Construct a table of valued of the function using the interval of -5 to 5.
j(t)=1/t
1
Expert's answer
2021-10-06T12:33:50-0400

Interval of t is -5 to 5

"At \space \space t=-5\\j(-5)=-\frac{1}{5}\\

At \space \space t=-4\\j(-4)=-\frac{1}{4}\\

At \space \space t=-3\\j(-3)=-\frac{1}{3}\\

At \space \space t=-2\\j(-2)=-\frac{1}{2}\\

At \space \space t=-1\\j(-1)=-1\\
At \space \space t=-0\\j(-0)=\frac{1}{0}=

 ∞"


"At \space \space t=1\\j(1)=1\\

At \space \space t=2\\j(2)=\frac{1}{2}\\

At \space \space t=3\\j(3)=\frac{1}{3}\\

At \space \space t=4\\j(4)=\frac{1}{4}\\

At \space \space t=5\\j(5)=\frac{1}{5}"